if a cos 2Φ + b sin 2Φ =c has p & q as its solutions, then prove that


tan p + tan q =2b/a+c      & tan p * tan q = c-a/a+c.

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear nishant

a cos 2Φ + b sin 2Φ =c

a(1-tan2Φ)/(1+tan2Φ)   +  b (2tanΦ)/(1+tan2Φ) =c

 tan2Φ(a+c)  +tanΦ(-2b) +(c-a)=0

now tan p  and tanq  are the solution of this equation

tanp + tanq = 2b/(a+c)

tanp*tanq = (c-a)/(c+a)

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin